### v3

@@ Abstract
Accurate depth and camera pose estimation is essential for achieving high-quality 3D visualisations in robotic-assisted surgery. Despite recent advancements in foundation model adaptation to monocular depth estimation of endoscopic scenes via self-supervised learning (SSL), no prior work has explored their use for pose estimation. These methods rely on low rank-based adaptation approaches, which constrain model updates to a low-rank space. We propose Endo-FASt3r, the first monocular SSL depth and pose estimation framework that uses foundation models for both tasks. We extend the Reloc3r relative pose estimation foundation model by designing Reloc3rX, introducing modifications necessary for convergence in SSL. We also present DoMoRA, a novel adaptation technique that enables higher-rank updates and faster convergence. Experiments on the SCARED dataset show that Endo-FASt3r achieves a substantial $10\%$ improvement in pose estimation and a $2\%$ improvement in depth estimation over prior work. Similar performance gains on the Hamlyn and StereoMIS datasets reinforce the generalisability of Endo-FASt3r across different datasets. Our code is available at: Endo-FASt3r.
@@ 2.1 Preliminaries
DoRA: Proposed to improve LoRA by drawing influence from Weight Normalization [20], DoRA decomposes the weight matrix into its magnitude and direction to accelerate optimisation [13]. The weight decomposition of $W\in\mathbb{R}^{d\times k}$ can be written as $W=m\frac{V}{||V||_{c}}$ where $m\in\mathbb{R}^{1\times k}$ is the magnitude vector, $V\in\mathbb{R}^{d\times k}$ is the directional matrix, with $||\cdot||_{c}$ being the column-wise vector norm of a matrix. At initialisation, given the pre-trained weight $W_{0}$, we set $m=||W_{0}||_{c}$ and $V=W_{0}$. LoRA with rank $r$ is then applied only to the directional component while the magnitude component is kept trainable. The updated weight $W^{\prime}$ is:
$W^{\prime}=\underline{m}\frac{V+\Delta V}{||V+\Delta V||_{c}}=\underline{m}\frac{W_{0}+\underline{BA}}{||W_{0}+\underline{BA}||_{c}}$ (4)
@@
Figure 1: EndoFAst3r SSL monocular depth and pose estimation pipeline.
@@ 2.2 Proposed Framework: Endo-FASt3r
DoMoRA: To benefit from the faster convergence of DoRA and the full-rank update space of MoRA, we introduce DoMoRA. Following Eqs. 43 and given the pre-trained weight $W_{0}^{d\times k}$, rank $r$, magnitude $m$, DoRA low-rank matrices $B^{d\times r}$,$A^{r\times k}$ and MoRA square matrix $M^{r\times r}$, the new layer with DoMoRA is:
$h(x)=(\underline{m}\frac{W_{0}+\underline{BA}}{||W_{0}+\underline{BA}||_{c}})x+f_{\text{decomp}}\left(Mf_{\text{comp}}\Big{(}x\right)\Big{)}$ (6)
In which $f_{\text{comp}}$ and $f_{\text{decomp}}$ follow the rotation and truncation methods[11]. Endo-FASt3r incorporates DoMoRA in the query and value matrices of all transformer blocks (Fig. 2). Inspired by the hierarchical feature learning nature of neural networks, we adopt the vector rank approach presented in [29]. The optimised rank vectors used are presented in the supplementary material.
@@ 3 Implementation details
Training: Endo-FASt3r utilises the Adam optimiser with a learning rate of 0.0001 decaying every 10 steps. Our model was trained on an A100 GPU for 10 epochs with a batch size of 4 for $20$ hours.
Datasets and Evaluation Protocol: Endo-FASt3r has been trained and evaluated on the SCARED dataset  [2], captured with a da Vinci Xi from porcine cadavers using structured light for depth collection. Following the split in [22], 15,351 frames were used for training, 1,705 for validation, and 551 for depth evaluation. Meanwhile pose evaluation utilised two trajectories of length 410 and 833 frames defined in [22]. To examine generalisability, we used the laparoscopic Hamlyn dataset following [19] for the depth module. For the pose module, we extracted frames 9,780–10,980 from sequence P3 of the StereoMIS dataset [9], which exhibit significant tissue deformation and camera motion in an in vivo setup. StereoMIS ground-truth poses were derived from the kinematics of the endoscope. The Depth evaluation metrics, mainly the absolute relative error (AbsRel), were adopted from [22]. Before evaluation, depth maps were scaled by median scaling [31]. The Absolute trajectory error (ATE) [14] was used for pose evaluation. More details are presented in the supplementary material.
@@ 4 Results
We primarily compare Endo-FASt3r against SOTA reprojection-based SSL approaches for monocular depth and pose estimation  [3, 8, 17, 28, 29] in RAS. To assess the generalisability of non-endoscopic foundation models DA V2 and Reloc3r, we also include their zero-shot performance. We use the reported metrics for all comparison methods in Table LABEL:restable [29, 6].
Table 1: Quantitive evaluation of Endo-FASt3r on the SCARED and Hamlyn datasets. ATE-T1/T2 denote ATE for trjactorries 1/2. Total/Train denote the total/trainable (millions) depth module parameters. Speed (ms) is reported for the depth module.
@@
Hamlyn Endo Depth & Motion [19] 0.185 5.424 16.100 0.732 - - - - -
@@
To ensure fairness, we excluded SSL methods with extra modules beyond depth and pose in their reprojection, such as camera intrinsics estimation or image intrinsic decomposition, which add extra self-supervision signals [6, 12]. Even if included, Endo-FASt3r still excels significantly in pose estimation and is comparable or superior in depth estimation, proving its high performance without requiring additional parameters.
Pose estimation: Evaluation on the rigid SCARED dataset shows Endo-FASt3r outperforming all SOTA methods, improving by $6.64\%$ and $12.04\%$ average $9.34\%$ over the second-best approach, DARES (with a CNN-based pose module), for trajectories 1 and 2 respectively (Table LABEL:restable). Fig. 3 presents this improvement on trajectory 2, where occlusion by surgical fluid is effectively addressed by Endo-FASt3r (please refer to the supplementary material). To assess generalisability to a non-rigid case with RAS-like tissue deformation and camera motion, we evaluated on the StereoMIS dataset, comparing against DARES, using its public code (Fig. 5). Endo-FASt3r outperforms DARES by $7.13\%$ (Table 3).
@@
We also present 3D reconstructions on the SCARED dataset in Fig. 5, which shows Endo-FASt3r better capturing darker areas with fewer artifacts. For further visual results of depth, pose, and 3D reconstruction, please refer to the supplementary material.
@@ 5 Conclusion
We present Endo-FASt3r, a self-supervised monocular depth and pose estimation framework for RAS. We introduce Reloc3rX, the first development of a foundation model for pose estimation in endoscopic scenes. We also introduce DoMoRA, a novel adaptation technique that enables both low-rank and full-rank updates while benefitting from faster convergence. As the first SSL-based method that uses foundation models for both depth and pose estimation, Endo-FASt3r achieves a substantial improvement ranging from $7\%-10\%$ in the ATE metric for pose estimation and $2\%$ in the AbsRel metric in depth estimation over prior methods, as shown across three public datasets. Future work will explore the use of video foundation models [26] to improve the temporal stability of our pipeline.

### v4

@@ Abstract
Accurate depth and camera pose estimation is essential for achieving high-quality 3D visualisations in robotic-assisted surgery. Despite recent advancements in foundation model adaptation to monocular depth estimation of endoscopic scenes via self-supervised learning (SSL), no prior work has explored their use for pose estimation. These methods rely on low rank-based adaptation approaches, which constrain model updates to a low-rank space. We propose Endo-FASt3r, the first monocular SSL depth and pose estimation framework that uses foundation models for both tasks. We extend the Reloc3r relative pose estimation foundation model by designing Reloc3rX, introducing modifications necessary for convergence in SSL. We also present DoMoRA, a novel adaptation technique that enables higher-rank updates and faster convergence. Experiments on the SCARED dataset show that Endo-FASt3r achieves a substantial $10\%$ improvement in pose estimation and a $2\%$ improvement in depth estimation over prior work. Similar performance gains on the Hamlyn and StereoMIS datasets reinforce the generalisability of Endo-FASt3r across different datasets. Our code is available at: https://github.com/Mona-ShZeinoddin/Endo_FASt3r.git.
@@ 2.1 Preliminaries
[…]
@@
Figure 1: EndoFAst3r SSL monocular depth and pose estimation pipeline based on the two foundation models, DA V2 and Reloc3rX.
@@ 2.2 Proposed Framework: Endo-FASt3r
DoMoRA: To benefit from the faster convergence of DoRA and the full-rank update space of MoRA, we introduce DoMoRA. Following Eqs. 43 and given the pre-trained weight $W_{0}^{d\times k}$, rank $r$, magnitude $m$, DoRA low-rank matrices $B^{d\times r}$$A^{r\times k}$ and MoRA square matrix $M^{r\times r}$, the new layer with DoMoRA is:
[…]
In which $f_{\text{comp}}$ and $f_{\text{decomp}}$ follow the rotation and truncation methods[11]. Endo-FASt3r incorporates DoMoRA in the query and value matrices of all transformer blocks (Fig. 2). Inspired by the hierarchical feature learning nature of neural networks, we adopt the vector rank approach presented in [29].
@@ 3 Implementation details
Training: Endo-FASt3r utilises the Adam optimiser with a learning rate of 1e41e-4 decaying every 10 steps. Training was performed on an A100 GPU for 10 epochs with a batch size of 4 for $20$ hours.
Datasets and Evaluation Protocol: Endo-FASt3r has been trained and evaluated on the SCARED dataset  [2], captured with a da Vinci Xi from porcine cadavers using structured light for depth collection. Following the split in [22], 15,351 frames were used for training, 1,705 for validation, and 551 for depth evaluation. Meanwhile pose evaluation utilised two trajectories of length 410 and 833 frames defined in [22]. To examine generalisability, we used the Hamlyn dataset following [19] for the depth module. For the pose module, we extracted frames 9,780–10,980 from sequence P3 of the StereoMIS dataset [9], which exhibit significant tissue deformation and camera motion in an in vivo setup. StereoMIS ground-truth poses were derived from the kinematics of the endoscope. The Depth evaluation metrics, mainly the absolute relative error (AbsRel), were adopted from [22]. Before evaluation, depth maps were scaled following [31]. The Absolute trajectory error (ATE) [14] was used for pose evaluation.
@@ 4 Results
We primarily compare Endo-FASt3r against SOTA reprojection-based SSL approaches for monocular depth and pose estimation  [3, 8, 17, 28, 29] in RAS. To assess the generalisability of non-endoscopic foundation models DA V2 and Reloc3r, we also include their zero-shot performance. We use the reported metrics for all comparison methods in Table LABEL:restable [29, 6]. To ensure fairness, we excluded SSL methods with extra modules beyond depth and pose in their reprojection, such as camera intrinsics estimation or image intrinsic decomposition, which add extra self-supervision signals [6, 12]. Even if included, Endo-FASt3r still excels significantly in pose estimation and is comparable or superior in depth estimation, proving its high performance without requiring additional parameters.
Table 1: Quantitive evaluation on the SCARED and Hamlyn datasets. ATE-T1/T2 denote ATE for trjactorries 1/2. Total/Train denote the total/trainable (millions) depth module parameters. Speed (ms) is reported for the depth module.
@@
Hamlyn
Endo Depth & Motion [19] 0.185 5.424 16.100 0.732 - - - - -
@@
Pose estimation: Evaluation on the rigid SCARED dataset shows Endo-FASt3r outperforming all SOTA methods, improving by $6.64\%$ and $12.04\%$ - average $9.34\%$ - over the second-best approach, DARES (with a CNN-based pose module), for trajectories 1 and 2 respectively (Table LABEL:restable). Fig. 3 presents this improvement on trajectory 2, where occlusion by surgical fluid is effectively addressed by Endo-FASt3r. To assess generalisability to a non-rigid case with RAS-like tissue deformation and camera motion, we evaluated on the StereoMIS dataset, comparing against DARES, using its public code (Fig. 5). Endo-FASt3r outperforms DARES by $7.13\%$ (Table 3). The pose module operates in real time, with a speed of $26ms$. Other works do not report this quantity.
@@
We also present 3D reconstructions on the SCARED dataset in Fig. 5, which shows Endo-FASt3r better capturing darker areas with fewer artifacts.
@@ 5 Conclusion
We present Endo-FASt3r, a self-supervised monocular depth and pose estimation framework for RAS. We introduce Reloc3rX, the first development of a foundation model for pose estimation in endoscopic scenes. We also introduce DoMoRA, a novel adaptation technique that enables both low-rank and full-rank updates while benefitting from faster convergence. As the first SSL-based method that uses foundation models for both depth and pose estimation, Endo-FASt3r achieves a substantial improvement ranging from $7\%-10\%$ in the ATE metric for pose estimation and $2\%$ in the AbsRel metric in depth estimation over prior methods, as shown across three public datasets. Future work will explore the use of video foundation models [26] to improve temporal stability.
{credits}
5.0.1 Acknowledgements
This work was supported in whole, or in part, by the the UKRI Centre for Doctoral Training (CDT) in AI-enabled Healthcare Systems [EP/S201612/1], the Department of Science, Innovation and Technology (DSIT), the Royal Academy of Engineering under the Chair in Emerging Technologies scheme, and the Engineering and Physical Sciences Research Council (EPSRC) under grants [EP/W00805X/1, EP/Z534754/1, UKRI145]. For the purpose of open access, the author has applied a CC BY public copyright licence to any author accepted manuscript version arising from this submission.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.